Solve initial-value problem with u(o)= -4
du dt 35 + 7u + 5t + ut
Answer #1

35+74+5ttut, U(o)=-4 du dt at du = (5+1) (747) acl au = (7+t) dt 5tu integrating en (5+4)= 7++ t² + c 2 2 te using Condition